OpenStudy (anonymous):

I'm having trouble integrating (100+2t)^3/2. I get [5(100+2t)^5/2]/5/2, which I think should be simplified to just 2(100+2t)^5/2. But the answer is just (100+2t)^5/2. What am I doing wrong?

ganeshie8 (ganeshie8):

you may use wolfram to verify : http://www.wolframalpha.com/input/?i=%5Cint+%28100%2B2t%29%5E%283%2F2%29+dt

OpenStudy (anonymous):

yeah wolfram alpha shows me the right answer but it doesn't show me how it got to that answer.

OpenStudy (anonymous):

I'm not sure what step I am not doing, or doing incorrectly.

ganeshie8 (ganeshie8):

\[\large \int (100+2t)^{\frac{3}{2}} ~dt\] substitute \(u = 100+2t\) \(du = 2dt \) \(dt = \dfrac{du}{2}\)

ganeshie8 (ganeshie8):

the integral becomes : \[\large \int u^{\frac{3}{2}} \frac{du}{2}\]

ganeshie8 (ganeshie8):

\[\large \frac{1}{2}\int u^{\frac{3}{2}}~dt\]

ganeshie8 (ganeshie8):

integrate ^^